# Weakly Supervised Learning for Facial Behavior Analysis: A Review

### v3

@@ 2.3 Inaccurate Supervision
In this category, dense labels are provided for the entire dataset as in the supervised learning scenario, yet the labels tend to be noisy. Since inaccurate labels degrade the performance of the prediction model, the goal is to overcome the challenges of noisy labels by training robust models that provide accurate predictions on test data. In the case of expression detection, labeling frames of the video with the corresponding expressions is a laborious task, and annotators are more vulnerable to mislabeling frames of the videos. For instance, the expression ”sadness” may look similar to the expression ”neutral”. It is therefore important to handle the problem of noisy expression labels and predict accurate expressions. The ambiguity is even more pronounced for compound expressions, with a combination of more than one expression. The problem of inaccurate annotations is more prevalent in the case of regression, where the intensity of expressions or AUs is estimated. Since more variation is possible for dimensional regression, annotations of dimensional regression tend to have more noise. The scenario of inaccurate annotations for expression recognition in videos is shown in Figure 3d, and state-of-the-art approaches for expression detection are discussed in Section 3.3.1. The framework of inaccurate annotations for expression recognition can also be extended to AU recognition, where AU labels are considered to be noisy. The framework of inaccurate annotations for AU recognition is shown in Figure 4d, and state-of-the-art approaches for AU detection are discussed in Section 3.3.2.
3 Methods for Classification
In this section, existing WSL methods specialized for classification are categorized into 3 types: inexact, incomplete, and inaccurate annotations, in the context of expression recognition and AU recognition applications. We also summarize experimental results obtained with these methods using common evaluation protocols and datasets.
3.1 Inexact Annotations
This subsection reviews WSL approaches related to inexact annotations, where annotations are provided at a global coarse level instead of dense low-level annotations.
3.1.1 Expression Recognition
In the case of inexact annotations, expression labels are provided at sequence-level and the goal of the task is to localize and predict the expressions of test sequences.
@@ 3.1.1 Expression Recognition
Instance Level Approaches: Expression recognition with MIL was initiated by Sikka et al. [28], where automatic pain localization was achieved using an instance-level classifier based on MILBOOST [29] and bag-level labels are predicted based on the maximum probabilities of the instances of the corresponding bag. Wu et al. [30] further enhanced this approach by incorporating a discriminative Hidden Markov Model (HMM) based instance level classifier with MIL instead of MILBOOST to efficiently capture the temporal dynamics. Chen et al. [31] explored pain classification with MIL using the relationship between AUs and pain expression, showing that clustering the co-occurrence of AUs provides better performance than individual AUs.
@@
Bag Level Approaches: Unlike the aforementioned approaches that rely on a single concept assumption, Adria et al. [32] proposed a multi-concept MIL framework, i.e., multiple expressions in a video for estimating high-level (bag-level) semantic labels of videos. A set of $k$ hyper-planes is modeled to discriminate $k$ concepts (facial expressions) in the instance space and the bag level representation is obtained using the probability of bag for each concept. Similarly, Sikka et al. [33] also modeled the temporal dynamics of facial expressions in videos using discriminative templates (neutral, onset, apex, or offset). The ordering of discriminative templates in a video sequence is associated with a cost function that captures the likelihood of the occurrence of different temporal orders. Unlike conventional MIL methods, Huang et al. [34] developed a novel framework for personal affect detection with minimal annotation (PADMA) for handling user-specific differences based on the association between key facial gestures and affect labels. In particular, if an instance occurs frequently in bags of a given class, but not in others, the instance has a strong association with the label.
3.1.2 Action Unit Recognition
@@
Since any expression can be characterized as a combination of AUs, many psychological studies have shown that there exists a strong association between expressions and AUs [35]. Due to the laborious task of obtaining AU annotations compared to expressions and the close relation between expressions and AUs (see Table 1), many researchers have explored the problem of AU detection in the framework of WSL, where expression labels are considered as weak coarse labels for AUs. Ruiz et al. [36] investigated the possibility of learning AU classifiers using expression labels by exploiting the relationship between expressions and AUs. Each input sample is mapped to an AU, which is in turn mapped to an expression, and AU classification is considered a hidden task. Each expression classifier is learned before training using an empirical study of the relationship between expressions and AUs [4], which is used to learn AU classifiers using gradient descent. Instead of using only the relationship between basic expressions and the corresponding AU probabilities, Wang et al.[37] leveraged expression-dependent and expression-independent AU probabilities without using extra large-scale expression-labeled facial images. First, the domain knowledge of expressions and AUs is summarized, based on which pseudo-AU labels are generated for each expression. Then a Restricted Boltzmann Machine is used to model prior joint AU distribution from the pseudo-AU labels (RBM-P). Using a similar approach, Peng et al. [38] explored adversarial training for AU recognition instead of maximizing the log-likelihood of the AU classifier with regard to the learned AU label prior. Inspired by generative adversarial networks (GAN), AU classifiers are learned by minimizing the differences between AU output distribution from AU classifiers and pseudo-AU label distribution derived from the summarized domain knowledge (RAN).
@@
Unlike the prior approaches, Wang et al. [39] modeled the relation between expressions and AU probabilities as inequalities instead of exact probabilities, i.e., higher probabilities of occurrence have higher rankings than those with lower probabilities. Similarly, Zhang et al. [40] comprehensively summarized the domain knowledge and exploited both expression-dependent and expression-independent AU probabilities to model the relationship among AU probabilities (LP-SM). Inspired by the idea of dual learning, Wang et al. [41] integrated the task of face synthesis along with AU recognition (WSDL), where the latter is considered as the main task and the former as an auxiliary task. Specifically, AU labels are predicted using AU classifiers learned from the domain knowledge, and a synthetic face is generated using predicted AU labels.
3.1.3 Discussion
@@
In the case of instance-level approaches, Sikka et al. [28] and Chen et al. [31] explored MILBOOST for instance-level classification, whereas Wu et al. [30] used an HMM-based classifier. Sikka et al. [28] and Wu et al. [30] explored SIFT-based and facial landmark-based features respectively, whereas Chen et al. [31] used feature representations based on the relationship between pain and AUs, which performs better than that of Sikka et al. [28] and Wu et al. [30]. For bag-level approaches, the bag-level representation has been obtained based on two strategies: statistics and the ordering of multiple expressions. The majority of bag-level approaches [32, 33] leverage the statistics of multiple expressions in the sequence. However, Sikka et al. [28] explored the ordering of expressions in the sequence and showed significant improvement in performance. Most of these approaches rely on traditional features such as Bag of Words (BoWs), geometric or SIFT-based features, and classical models due to the limited amount of representative data. For AU classification with expression labels, most approaches used features based on facial landmark-based [37, 38, 41, 39], while Ruiz et al. [36] and Zhang et al. [40] used SIFT and LBP features, respectively. Most of the approaches used linear classifiers to obtain the probabilities of predicted AUs, whereas Ruiz et al. [36] used multi-class SVM. Among these approaches, various aspects of the relationships between expressions and AUs are explored. Ruiz et al. [36] explored the basic relationships between AUs and expressions. Wang et al. [37] and Peng et al. [38] improved the performance by analyzing the expression-dependent and expression-independent AUs. [39] and [40] explored the relationships in terms of relative probabilities using inequality constraints instead of exact probabilities between expressions and AUs. Unlike prior approaches, [41] explored AU recognition as a dual learning problem where AU recognition is achieved by incorporating face synthesis. Approaches for AU detection with weak expression labels have also been extended to the problem of incomplete AU annotations by deploying additional loss components for the partial AU annotations [36, 37, 38, 41, 39, 40].
3.2 Incomplete Annotations
3.2.1 Expression Recognition
@@
Pseudo-Labeling: The idea of pseudo-labeling is to predict the labels of unlabeled data, where the predicted pseudo-labels of higher confidence are in turn used to train the unlabeled data in a supervised setting. Happy et al. [42] explored expression classification even with low intensities. Initially, they train a CNN model with limited labeled data until adequate performance is achieved. Subsequently, model parameters are further updated by fine-tuning using a portion of unlabeled data with high-confidence predictions, obtained by the current model in every epoch. Li et al. [43] proposed an entropy minimization method based on adaptive confidence margin (Ada-CM) to obtain more reliable pseudo-labeled labels, where they used high-confidence unlabeled samples to obtain pseudo-labeled labels and samples with low confidence for contrastive learning to obtain feature representations.
@@
Consistency Regularization: It is based on the assumption that realistic perturbations of the same input generate consistent predictions. Jiang et al. [44] addressed the problem of limited representative training data by exploiting the large-scale unlabeled data in a teacher-student framework based on a cross-guidance mechanism (Progressive Teacher). The teacher network enforces stronger learning ability and guides the student model, which is used to optimize the classification loss.
@@
Hybrid Methods: Some of the approaches have explored both pseudo-labeling and consistency regularization in a hybrid framework. Florea et al. [45] (margin-mix) improved the idea of center loss [46] by maximizing the distance between the centroids of the different classes in the loss function. Pseudo labels are estimated based on the distances to centroids of different classes and used along with mix-up augmentation [47] to avoid over-fitting. Fang et al. [48] proposed a novel framework of pseudo-labeling strategy (Rethink PL) by addressing the problem of class distribution mismatch between the labeled and unlabeled datasets.
@@
Unlike the above-mentioned approaches, Kurup et al. [49] explored Deep Belief Networks (DBN) to train on unlabeled data, which is further fine-tuned on labeled data.
3.2.2 Action Unit Recognition
@@
Missing Labels: Song et al. [50] developed a Bayesian framework by encoding the sparsity. Only very few AUs are active at any moment, and the co-occurrence structure of AUs, i.e., overlapping of AUs in multiple groups via compressed sensing and group-wise sparsity inducing priors (BGCS). Wu et al. [51] proposed a multi-label learning framework with missing labels (MLML) by enforcing the constraints of consistency between predicted labels and provided labels (label consistency) as well as with label smoothness, i.e., labels of similar features should be close to each other along with modeling the co-occurrence relationships among AUs. However, Yongqiang et al. [52] found that the constraint of label smoothness with shared feature space among AUs is violated for the task of AU recognition due to the diverse nature of the occurrence of AUs. That is, different AUs occur in different face regions, thereby features selected for one AU classifier are not discriminative for other AU classifiers. Discriminative features are therefore learned for each AU class before deploying the constraint of label smoothness. Li et al. [53] further extended the idea of [52] to address the problem of class imbalance in two aspects – the number of positive AUs being much smaller than the negative AUs in each sample (image) and the rate of positive samples of different AUs being significantly different.
@@
Incomplete Labels: Wang et al. [54] deal with incomplete AU labels but complete expression labels by modeling the dependencies among AUs and the relationship between expressions and AUs with Bayesian network (BN) using maximum likelihood estimation. Peng et al. [55] explored an adversarial GAN-based approach with dual learning by leveraging domain knowledge of expressions and AUs along with facial image synthesis from predicted AUs. Specifically, the probabilistic duality between tasks and the dependencies among facial features, AUs, and expressions are explored in an adversarial learning framework.
@@
3.2.3 Discussion
For expression classification, approaches based on pseudo-labeling [42] have shown promising performance, however, incorrect pseudo-labels may degrade the performance. Therefore, [43] improved the reliability of pseudo-labels based on adaptive confidence margin. Recently, hybrid methods [45, 48] have leveraged both strategies to improve performance. DL models have also been explored to leverage large unlabeled data to deal with limited annotations [42, 43, 45, 48]. Most of these methods explore Resnet-18 [44], WideResnet-28-2 [45] or both [43, 48]. For AU detection, two research directions have been explored: missing and incomplete labels. Most approaches for AU detection with missing and incomplete labels used features based on facial landmark-based features [52, 51, 53, 54, 56], whereas Song et al. [50] used HoG features, and [56, 57] used deep models. [50] and [54] used Bayesian classifiers for AU recognition, while [56] and [55] used multi-class SVM and GANs, respectively, for classification. Most approaches to missing labels explored the property of co-occurrence of AUs [50, 51, 52] to deal with missing AUs. For incomplete AU labels, Wang et al. [54] and Peng et al. [55] leveraged the expression labels along with incomplete AU labels, while Wu et al. [56] and Niu et al. [57] rely only on incomplete AU labels.
3.3 Inaccurate Annotations

### v4

@@ II-C Inaccurate Supervision
In this category, dense labels are provided for the entire dataset similar to that of supervised learning, but the labels tend to be noisy. Since inaccurate labels degrade the system performance, the goal is to overcome the challenges of noisy labels by training robust models that can provide accurate predictions on test data. One of the major strategies to deal with inaccurate labels is crowd-sourcing [29], where the labels are obtained from multiple annotators. In the case of expression detection, labeling frames of the video is a laborious task, thereby annotators are more vulnerable to mislabeling frames of the videos. For instance, the expression ”sadness” may look similar to the expression ”neutral”. The ambiguity is even more pronounced for compound expressions, with a combination of more than one expression. The problem of inaccurate annotations is more prevalent in the case of regression, where the intensity of expressions or AUs is estimated. Since more variation is possible for dimensional regression, annotations of dimensional regression tend to be highly noisy. The framework of inaccurate annotations for expression recognition can also be extended to AU recognition, where AU labels are considered to be noisy. The framework of inaccurate annotations for expressions and AUs are shown in Figure 3d and Figure 4d respectively.
III Methods for Classification
In this section, WSL methods for both expression and AU recognition are presented each WSL scenario along with the critical analysis, emphasizing the strengths and limitations. The comparison of the experimental results is also provided based on common evaluation protocols and datasets.
III-A Inexact Annotations
III-A1 Expression Recognition
@@
.                                                      Instance Level Approaches: Expression recognition with MIL was initiated by Sikka et al. [30], where automatic pain localization was achieved using an instance-level classifier based on MILBOOST [31] and bag-level labels are predicted based on the maximum probabilities of the instances of the corresponding bag. Wu et al. [32] further enhanced this approach by incorporating a discriminative Hidden Markov Model (HMM) for instance level classifier with MIL instead of MILBOOST to efficiently capture the temporal dynamics. Chen et al. [33] explored pain classification with MIL using the relationships between AUs and pain expression, underscoring that clustering the co-occurrence of AUs provides better performance than individual AUs.
@@
Bag Level Approaches: Unlike the aforementioned approaches that rely on a single concept assumption, Ruiz et al. [34] proposed a multi-concept MIL framework, i.e., multiple expressions in a video for estimating high-level (bag-level) semantic labels of videos. A set of $k$ hyper-planes is modeled to discriminate $k$ concepts (facial expressions) in the instance space and the bag level representation is obtained using the probability of bag for each concept. Similarly, Sikka et al. [35] also modeled the temporal dynamics of facial expressions in videos using discriminative templates (neutral, onset, apex, or offset). The ordering of discriminative templates in a video sequence is associated with a cost function that captures the likelihood of the occurrence of different temporal orders. Unlike conventional MIL methods, Huang et al. [36] developed a novel framework for personal affect detection with minimal annotation (PADMA) for handling user-specific differences based on the association between key facial gestures and affect labels. In particular, if an instance occurs frequently in bags of a given class, but not in others, the instance has a strong association with the label.
III-A2 Action Unit Recognition
@@
Since any expression can be characterized as a combination of AUs, many psychological studies have shown that there exists a strong association between expressions and AUs [37]. Due to the laborious task of obtaining AU annotations compared to expressions and the close relation between expressions and AUs (see Table 1), many researchers have explored the problem of AU detection in the framework of WSL, where expression labels are considered as weak coarse labels for AUs. One of the early works was done by Ruiz et al. [38], where each input sample is mapped to an AU, which is in turn mapped to an expression, and AU classification is considered a hidden task. Each expression classifier is learned before training using an empirical study of the relationships between expressions and AUs [4], which is used to learn AU classifiers using gradient descent. Instead of using only the relationships between basic expressions and the corresponding AU probabilities, Wang et al.[39] leveraged expression-dependent and expression-independent AU probabilities without using extra large-scale expression-labeled facial images. First, the domain knowledge of expressions and AUs is summarized, based on which pseudo-AU labels are generated for each expression. Then a Restricted Boltzmann Machine is used to model prior joint AU distribution from the pseudo-AU labels (RBM-P). Using a similar approach, Peng et al. [40] explored adversarial training for AU recognition instead of maximizing the log-likelihood of the AU classifier with regard to the learned AU label prior. Inspired by generative adversarial networks (GAN), AU classifiers are learned by minimizing the differences between AU output distribution from AU classifiers and pseudo-AU label distribution derived from the summarized domain knowledge (RAN).
@@
Unlike the prior approaches, Wang et al. [41] modeled the relation between expressions and AU probabilities as inequalities instead of exact probabilities, i.e., higher probabilities of occurrence have higher rankings than those with lower probabilities. Similarly, Zhang et al. [42] comprehensively summarized the domain knowledge and exploited both expression-dependent and expression-independent AU probabilities to model the relationships among AU probabilities (LP-SM). Inspired by the idea of dual learning, Wang et al. [43] integrated the task of face synthesis along with AU recognition (WSDL), where the latter is considered as the main task and the former as an auxiliary task. Specifically, AU labels are predicted using AU classifiers learned from the domain knowledge, and a synthetic face is generated using predicted AU labels.
III-A3 Discussion
In the case of instance-level approaches, Wu et al. [32] showed that effectively capturing the temporal dynamics of expressions in videos plays a pivotal role for better performance. Though the relationships between expressions and AUs has been well-explored for AU classification [39, 40], they are not much explored for expression classification. Chen et al. [33] showed that leveraging the relationships between AUs and expressions can improve the performance of instance-level expression classification. Though initial approaches [32, 33, 30] using MIL are based on single-concept assumption, multi-concept approaches are more relevant for real-world scenarios. Effective modeling of the temporal dynamics of multiple concepts (expressions) in the videos plays a crucial role for bag-level classification. Ruiz et al. [34] explored probabilistic modeling of the multiple expressions in the videos, whereas [35] captured the temporal dynamics of expressions based on statistics (frequency) of expressions in videos. Unlike these two approaches, Sikka et al. [30] showed that modeling the temporal ordering of expressions across the frames of the video seems to be quite promising to effectively capture the temporal dynamics of expressions.
@@ III-A3 Discussion
Among AU classification approaches based on expression labels, various aspects of the relationships between expressions and AUs are explored. Ruiz et al. [38] explored the basic relationships between AUs and expressions. Wang et al. [39] and Peng et al. [40] improved the performance by analyzing the expression-dependent and expression-independent AUs. Compared to [39], [40] demonstrated that adversarial training using GANs can help in learning the hidden AU classifiers better than RBMs. Unlike these approaches, [41] and [42] explored the relationships in terms of relative probabilities using inequality constraints and showed that they can perform better than that of approaches relying on exact probabilities between expressions and AUs. In addition to the learning hidden AU classifiers from expression labels, [43] explored a novel direction of dual learning framework where the task of face synthesis is also deployed, which implicitly guides in learning better AU classifiers, achieving better performance than prior approaches. Approaches for AU detection with weak expression labels have also been extended to the problem of incomplete AU annotations by deploying additional loss components for the partial AU annotations [38, 39, 40, 43, 41, 42].
III-B Incomplete Annotations
III-B1 Expression Recognition
@@
. Pseudo-Labeling: Pseudo-labeling seeks to predict the labels of unlabeled data, where the predicted pseudo-labels of higher confidence are in turn used to train the unlabeled data in a supervised setting. Happy et al. [44] explored expression classification even with low intensities. Initially, they train a CNN model with limited labeled data until adequate performance is achieved. Subsequently, model parameters are further updated by fine-tuning using a portion of unlabeled data with high-confidence predictions, obtained by the current model in every epoch. Li et al. [45] proposed an entropy minimization method based on adaptive confidence margin (Ada-CM) to obtain more reliable pseudo labels, where they used high-confidence unlabeled samples to obtain pseudo labels, while exploiting the samples with low confidence for contrastive learning to improve the feature representations.
@@
Consistency Regularization: It is based on the assumption that realistic perturbations of the same input generate consistent predictions. Jiang et al. [46] explored two pairs of teacher-student models, where the outputs of teacher model is forced to be consistent with that of student model in a cross-guidance mechanism. The teacher enforces stronger learning ability and guides the student model, gradually achieving similar learning ability (Progressive Teacher).
@@
Hybrid Methods: Some of the approaches have explored both pseudo-labeling and consistency regularization in a hybrid framework. Florea et al. [47] (margin-mix) improved the idea of center loss [48] by maximizing the distance between the centroids of the different classes in the loss function. Pseudo labels are estimated based on the distances to centroids of different classes and used along with mix-up augmentation [49] to avoid over-fitting. Fang et al. [50] proposed a novel framework of pseudo-labeling strategy (Rethink PL) by addressing the problem of class distribution mismatch between the labeled and unlabeled datasets using clustering-based approach.
@@
Unlike the above-mentioned approaches, Kurup et al. [51] explored Deep Belief Networks (DBN) to train the selected features of unlabeled data, which is further fine-tuned on labeled data.
III-B2 Action Unit Recognition
@@
. Missing Labels: Song et al. [52] developed a Bayesian graphical framework (BGCS) to encode sparsity and co-occurrence using compressed sensing and group-wise sparsity inducing priors based on two key observations: only very few AUs are active at any moment (sparsity), and overlapping of AUs in multiple groups (co-occurrence). Wu et al. [53] proposed a multi-label learning framework with missing labels (MLML) by enforcing the constraints of consistency between predicted labels and provided labels (label consistency) as well as with label smoothness, i.e., labels of similar features should be close to each other along with modeling the co-occurrence relationships among AUs. However, Li et al. [54] found that the constraint of label smoothness with shared feature space among AUs is violated for the task of AU recognition due to the diverse nature of the occurrence of AUs. That is, different AUs occur in different face regions, thereby features selected for one AU classifier may not be discriminative for other AU classifiers. Discriminative features are therefore learned for each AU class before deploying the constraint of label smoothness. Li et al. [55] further extended the idea of [54] by addressing the problem of class imbalance in two aspects – the number of positive AUs being much smaller than the negative AUs in each sample (image) and the rate of positive samples of different AUs being significantly different.
@@
Incomplete Labels: Wang et al. [56] dealt with incomplete AU labels but complete expression labels by modeling the dependencies among AUs and the relationships between expressions and AUs with Bayesian network (BN) using maximum likelihood estimation. Peng et al. [57] explored an adversarial GAN-based approach with dual learning by leveraging domain knowledge of expressions and AUs along with facial image synthesis from predicted AUs. Specifically, the probabilistic duality between tasks and the dependencies among facial features, AUs, and expressions are explored in an adversarial learning framework.
@@
III-B3 Discussion
For expression classification, approaches based on pseudo-labeling [44] have shown promising performance, however, incorrect pseudo-labels may degrade the performance. Therefore, most of the prior approaches of pseudo-labeling rely only on high-confidence samples, ignoring the low-confidence samples. However, [45] showed that low-confidence samples can also be exploited to improve the feature representations along with high-confidence samples. It has been shown that exploiting the large-scale data without labels or with noisy labels seems to be promising line of research to deal with data-hungry DL models [44, 45, 46]. Recently, hybrid methods [47, 50] have leveraged both strategies of consistency regularization [46] and pseudo-labeling [44, 45], thereby achieving better performance. [47] explored center-loss for expression classification, which seems to be an interesting idea for estimating pseudo-labels, while still leveraging the mix-up augmentation to deal with over-fitting issues. In most of these approaches, the class distribution mismatch between the reliable labeled data and large-scale unlabeled data is often overlooked, which has been addressed by [50]. Majority of these methods explore Resnet-18 [46], WideResnet-28-2 [47] or both [45, 50].
For AU detection, two research directions have been explored: missing and incomplete labels. [52, 53, 54, 55] addressed the problem of missing labels, whereas [56, 57, 58, 59] focused on handling incomplete labels. The core idea of dealing with missing labels is to exploit the co-occurrence structure of AUs to estimate the missing labels. In addition to the co-occurrence structure, [52] leveraged the sparse nature of AUs whereas [53, 54, 55] exploited the constraints of label-consistency and label-smoothness. Compared to [53] and [54], [55] further showed improvement by addressing the problem of class imbalance among AUs along with these constraints. For incomplete AU labels, Wang et al. [56] and Peng et al. [57] leveraged the relationships between the expression labels and AUs along with incomplete AU labels, while Wu et al. [58] and Niu et al. [59] depend only on incomplete AU labels. [57] explored adversarial training for better modeling of the relationships between expressions and AUs similar to their previous work of AU recognition from expressions [40], whereas [58] used Bayesian networks. Along with adversarial training, [57] also leveraged dual learning to improve the modeling of relationships between expressions and AUs similar to that of [43]. The idea of leveraging dual learning seems to be promising by implicitly enforcing better learning of the hidden AU classifiers. [58] focused on estimating the AUs of unlabeled samples using prior relationships among AUs (co-occurrence), which is modeled using RBMs whereas [59] showed better improvement using co-regularization based on the idea of co-training.
III-C Inaccurate Annotations